Solve for x: log4 x-log4 (2x-1)= 1
Answer by jim_thompson5910(28715)

log4 x-log4 (2x-1)= 1


log4[x/(2x-1)]= 1


x/(2x-1)= 4^1


x/(2x-1)= 4


x= 4(2x-1)


x=8x-4


-7x=-4


x=-4/(-7)


x=4/7